What volume of 25% Na2CO3 solution will be required to neutralize 2.5L of 0.25N H2SO4?
1
Expert's answer
2020-11-02T08:56:39-0500

Na2SO3 (aq) + H2SO4 (aq) → SO2 (aq) + Na2SO4 (aq) + H2O (l)

0.25N H2SO4 = 0.125 M H2SO4; 2.5L 0.125M H2SO4 = 0,3125 mol H2SO4; So, we need 0.3125 mol Na2SO3; M(Na2SO3)=126 g/mol; m(Na2SO3)=126*0.3125=39,375g

39,375g=25%

100%=39,375*4=157.5g what equal to 157.5 ml of solution.